OpenStudy (anonymous):

Locate a point C on the x-axis such that AC+BC is minimized. C=(_,0)

OpenStudy (anonymous):

The endpoints of AB are A(0,2) and B(9,4)

jimthompson5910 (jim_thompson5910):

check out this page http://www.analyzemath.com/calculus/Problems/minimum_distance.html

jimthompson5910 (jim_thompson5910):

I recommend you read over where it says " METHOD 2: We "construct a virtual house" (projection) H on the other side of the river"

jimthompson5910 (jim_thompson5910):

let me know if that helps or not

OpenStudy (anonymous):

That just confuses me..

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

First plot the points A and B

jimthompson5910 (jim_thompson5910):

If we use method 2, we reflect point B over the x axis to get point B'

jimthompson5910 (jim_thompson5910):

Then you draw a straight line from A to B'

jimthompson5910 (jim_thompson5910):

The line we just drew goes through (3,0) so that is the point that minimizes the distance AC + BC this is because the shortest distance between any two points is a straight line and because CB = CB' (you can prove this using CPCTC and the hypotenuse leg postulate)

OpenStudy (anonymous):

Okay, that makes sense. is there an equation I can use to set it up to show my work?

OpenStudy (anonymous):

nice

jimthompson5910 (jim_thompson5910):

let C be the point (x,0) the distance from A(0,2) to C(x,0) can be calculated using the distance formula d = sqrt( (x1-x2)^2 + (y1-y2)^2 ) d = sqrt( (0-x)^2 + (2-0)^2 ) d = sqrt( x^2 + 4 ) so that is the distance from A to C

jimthompson5910 (jim_thompson5910):

the distance from C to B is d = sqrt( (x1-x2)^2 + (y1-y2)^2 ) d = sqrt( (x-9)^2 + (0-4)^2 ) d = sqrt( x^2 - 18x + 81 + 16 ) d = sqrt( x^2 - 18x + 97 )

jimthompson5910 (jim_thompson5910):

The total distance AC + BC is the sum of those two distances sqrt( x^2 + 4 ) + sqrt( x^2 - 18x + 97 ) you would then use calculus or a graphing calculator to find the min of this function to get the min distance

OpenStudy (anonymous):

Okay, thanks so much for your help. Could you help me with one more problem?

jimthompson5910 (jim_thompson5910):

sure, go ahead

OpenStudy (anonymous):

Fill in the values to complete the reflection of AB with matrix multiplication.

jimthompson5910 (jim_thompson5910):

That's all it says?

OpenStudy (anonymous):

Yeah

jimthompson5910 (jim_thompson5910):

well I found this page http://www.mathplanet.com/education/geometry/transformations/transformation-using-matrices but the problem is that I don't know what line we are reflecting over

jimthompson5910 (jim_thompson5910):

are we reflecting over the x axis? y axis? or some other line?

OpenStudy (anonymous):

x-axis

jimthompson5910 (jim_thompson5910):

what are the coordinates of A and B?

jimthompson5910 (jim_thompson5910):

same from that previous problem?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

On this page http://www.mathplanet.com/education/geometry/transformations/transformation-using-matrices it says that the reflection matrix for reflecting over the axis is this